(1 point) Book Problem 9 6x - 5 Suppose that f(x) x +7 (A) Use interval notation to indicate where f(x) is defined. If it isEach letter is a piece of one question. Therefore, fill in all of the blanks please. Thank you
Answer #1

f(x) = 62-5 x+7 + t () 47. (x+7)2 94 f (X) = (x+7)3 f (x) is defined when X+770 treo X7-7 is Domain: --00,-7) U (-7,00) (An( 0 ) Now, f60 is increasing it f (x) yo is f(X) = 47 since, (x+7) >0 ton allx (x+7) i f(x) > 0 for all x But at a=-7, t eN(6) $() is con cave up if f(x) > 0 down f (2) LO Foto con cave 94 >0 (x+7)3 -94 20 (x+7)3 94 Lo + (x+7)3 LO ² x+7 LO # x<-7